Montserrat Volcano Observatory, Montserrat, West Indies

Scientific Report 40
26 October 1996


The October 1 dome continues to grow and is now visible from some parts to the north of the volcano. Volcano-tectonic earthquake swarms, which had been a dominant feature of the seismicity prior to the 17/18 September eruption have returned and four such swarms were recorded during the week. The sulphur dioxide flux has returned to about 200-400 tonnes/day after a maximum of 1510 tonnes/day on 24 October. Changes in line lengths measured by the MVO GPS and Total Station have continued to show similar trends to those observed during the previous weeks.

Visual observations

No clear views of the dome were obtained at the beginning of the week, but a rockfall at 11:35 on 20 October was large enough to be heard from Harris and Perches Mountain. Any ash cloud which may have been generated by this event was obscured by very low cloud cover. Observations from Whites on 21 October indicated that there were no significant visible changes in the appearance of the dome. Very few rockfalls were observed on the eastern face of the October 1 dome.

By 22 October, the height of the dome had increased from 808 m (2640 ft) on October 18 to 829 m (2710 ft). Several rockfalls were observed on the eastern face, and the dome appeared to be more active than in previous days. Rockfalls were occurring from the northern side of the scar and vigorous steaming was observed from the south of the October 1 dome.

Good visibility on 23 October allowed observations of some moderate-sized rockfalls from the October 1 dome which produced small ash clouds. Collapse of a small spine was observed and new growth was seen on the south side of the eastern face of the dome (FIGURE 1.). The new dome became visible from Farrells on the northern flank of the volcano on 23 October.

Viewing conditions deteriorated on 24, 25 and 26 October and no observations were possible. One brief view of the October 1 dome indicated that no major changes had occurred although the dome appeared to be slightly bigger than when it was last sighted.


Earthquake types: 20 to 26 October 1996

These earthquake counts are of events that triggered the short-period seismic network event recording system between 0000 to 2400 each day.

Volcano-tectonic earthquakes and rockfall signals have dominated the seismicity. The general level of activity increased during the latter part of the week with the re-emergence of volcano-tectonic earthquake swarms. The number of rockfalls has decreased further although their size has continued to increase.

Date       VT        LP        Hybrid    RF        Tremor
20 OCT 96  8         2         0         7         none
21 OCT 96  36        1         0         7         low
22 OCT 96  5         2         0         12        low
23 OCT 96  13        2         0         12        moderate
24 OCT 96  16        2         0         6         moderate
25 OCT 96  18        0         0         6         low
26 OCT 96  80        1         0         3         low

After a brief reduction in activity on the first day of the week, volcano-tectonic earthquakes increased during the remainder of the week occurring in swarms of variable duration. Four earthquake swarms were recorded during the week; 08:33-13:10 on 21 October (36 events); 20:47 on 23 October to 03:30 on 24 October (25 events); 00:08-06:44 on 25 October (18 events) and 03:11-21:18 on 26 October (80 events). Most of the earthquakes were located at shallow depths beneath the crater. The VT swarms are quite similar to swarms recorded in August and September and are thought to be related to magma movements at shallow depths beneath the volcano.

A few volcano-tectonic earthquakes were recorded outside of the swarm periods. One on 22 October was located at a depth of 5 km to the north of the crater. Two other events were located at about 2 km depth beneath St Georges Hill.

The number of rockfall signals remained low, with a further reduction in numbers towards the end of the week. The larger events produced small ash clouds which caused ashfall on the western flank of the volcano.

Three regional earthquakes were recorded during the period. One of these which occurred on 24 October was located to the east of Barbuda.

Very few long-period earthquakes have been recorded. Two rockfall signals recorded on 20 October had sharp, long-period starts that are associated with small explosions, such events have been absent since then.

The new broadband network is now fully operational. Analysis of the tremor recorded in the Gages area by the broadband sensor shows that it is narrow band and high-frequency, dominated by frequencies of 4 to 30 Hz. This tremor is probably due to a local source, possibly steam movement at Gages Soufriere.

Ground Deformation

The MVO Total Station was used to measure line lengths on all except the southern triangle during the week. Low cloud cover severely limited visibility at the volcano during the latter part of the week so that no measurements were possible from the 24 October onwards.

Measurements were carried out on the northern triangle between Windy Hill, Farrells and St Georges Hill on 21 October. The results indicate that line lengths between St George's Hill and Farrells lengthened by 3 mm, while the Windy Hill to Farrells line lengthened by 5 mm. The accuracies of these results indicate that no significant changes in the line lengths have been detected.

The eastern triangle was measured on 22 and 23 October. On 22 October, line lengths shortened by 3-4 mm since the previous measurement on 19 October. Measurement of this triangle on the 23 October indicated no changes in line lengths since the previous day.

The Chances Peak to Amersham lines were measured on 22 October. Line lengths between Amersham and Chances Peak steps shortened by 7 mm since these were last measured on 26 September. These continue to show essentially no significant change.

A minor glitch in the EDM database was discovered during the week. The reported lengthening of 3.8 cm on the Galways to Chances Peak line was in fact a shortening of 1.8 mm over the period 06 - 19 October.

GPS surveys were conducted on various networks around the volcano during the week. The Northnet which includes the northern part to the island was occupied on 20 October; Bignet which straddles the volcano was occupied on 22 October and the Eastnet which consists of locations on the eastern flank of the volcano was completed on 25 October. The results do not indicate significant change at the 95% confidence level.

Construction of a new GPS site for UPRM program of permanent GPS monitoring began at Whites during the latter half of the week. This site should be completed during the next week and would replace the Hermitage site which was severely damaged by ballistic projectiles from the 17/18 September eruption.

Dome volume measurements

During the excellent weather on 23 October another kinematic survey of the dome was completed using the range-finding binoculars and GPS equipment. The results of this survey along with all surveys completed thus far on the October 1 dome are presented below. The effusion rate for the dome between the surveys has been estimated to be 1.8 m3s-1.

Date            Dome Volume
02 OCT 96       2.0 x 105 m3
03 OCT 96       4.7 x 105 m3
10 OCT 96       1 x 106 m3
18 OCT 96       3 x 106 m3 (corrected from 4.7 106 m3 since 
				the preliminary results in 
				Scientific Report #39)
23 OCT 96       3.7 x 106 m3

The dome material continues to look significantly darker than the pre-September 17th dome. No sample has, as yet, been obtained.

Gas, Ash and Rainwater Measurements

Determination of sulphur dioxide flux in the volcanic plume was done for most of the week although measurements were hampered by adverse weather conditions at the beginning of the week. COSPEC measurements are normally done by mounting the instrument in an automobile and running traverses beneath the plume along the west coast road. Strong northerly winds at the beginning of the week took the plume over the central and more inaccessible parts of the island.

COSPEC measurements conducted along the west coast road from Cork Hill to Gingoes on 21, 23, 24 and 26 October. The average flux readings were 580, 449, 1519 and 290 tonnes per day respectively. The value obtained on 24 October is the highest recorded since the beginning of the volcanic crisis. This relatively high value is thought to be due to strong degassing of magma at shallow depths beneath the volcano.

The SO2 tubes at sites around Amersham and Plymouth are collected every two weeks and sent to the UK for analysis. No results have been obtained this week. No ash was sampled this week. Weekly collecting and analysis of rainwater continues.

Staff Changes
Rod Stewart, BGS

Prof Geoff Wadge, Reading University, U.K.

Montserrat Volcano Observatory